DATA GIVEN FOR THE OLD LOAN AND THE NEW LOAN

RUSH RUSH FINANCE ....PERSONAL Marginal tax rate 6% Resale Plan (Months) 60 MORTGAGE TERMS Original Mortgage Amount $150,000.00 Current Mortgage Rate 5.75% Original Term (Years) 20 Months Paid 117 New Mortage Rate 3.00% New Term (YEARS) 15 Points 0 REFINACING FEES Application $200.00 Title $2,000.00 Legal $200.00 Other $10,779.85 Points 0 Total Fees $13,179.85 ANALYSIS CURRENT PROPOSED Mortgage Amount 97, 703.26 $97,703.26 ANALYSIS Mortgage Payment Months to Recover Refinancing cost Mortage Balance at Resale (Valor Presente del MA - c26 Principal Repaid to Resale ( - MA - MBAR = c26 c29) Total Payments to Resale Total Interest to Resale Tax Deduction on Interest Net Interest Costo to Resale Interest Savings (COSTS) TOTAL SAVINGS COST (COSTS)

COURSE : FINANCE: INCLUDED DATA FOR OLD LOAM AND NEW LOAN TO RECALCULATE MORTGAGE PAYMENTS ETC

PERSONAL
MARGINAL TAX RATE 6%
RESALE PLAN (MONTHS) 60
MORTGAGE TERMS
ORIGINAL MORTGAGE AMOUNT $150,000.00
CURRENT MORTGAGE RATE 5.75%
ORIGINAL TERM (YEARS) 20
MONTHS PAID 117
NEW MORTGAGE RATE 3.00%
NEW TERM (YEARS) 15
POINTS 0
REFINANCING FEES
APPLICATION $200.00
TITLE $2,000.00
LEGAL $200.00
OTHER $10,779.85
POINTS 0
tOTAL FEES $13,179.85
ANALYSIS CURRENT PRPOSED
NORTGAGE AMOUNT $97,703.26 $97,703.26
   ANALYSIS FOR THE NEW LOAN : FIND
MORTGAGE PAYMENT
MONTHS TO RECOVER REFINANCING COST
MORTGAGE BALANCE AT RESALE (PRESENT VALUE)
PRINCIPAL REPAID TO RESALE
TOTAL PAYMENTS TO RESALE
TOTAL INTEREST TO RESALE
TAX DEDUCTION ON INTEREST
NET INTEREST COST TO RESALE
INTEREST SAVINGS (COSTS)
TOTAL SAVINGS COST
Answer #1

Soln :

ANALYSIS For New loan
MORTAGE AMOUNT 97703.26
MORTGAGE PAYMENT 674.72
MONTHS TO RECOVER REFINANCING COST 180
MORTGAGE BALANCE AT RESALE (PRESENT VALUE) 69865
PRINCIPAL REPAID TO RESALE 27838.26
TOTAL PAYMENTS TO RESALE 40483.2
TOTAL INTEREST TO RESALE 12644.94
TAX DEDUCTION ON INTEREST 758.70
NET INTEREST COST TO RESALE 11886.24
INTEREST SAVINGS (COST) 10749.76
TOTAL SAVINGS COST 13179.85

Please refer the table, here we have calculated mortgage payment for new loan a sPMT (3%/12,180,-97703.26,0) = 674.72

Total payments to resale that is at 60 months = 674.72*60

Also, we have calculated the interest cost for 60 months for old loan adding interest of each month .

Use Excel to calculate PMT, principal paid, interest and principal remaining, you will get these values for both old and new loan

Interest of previous loan = 22636 and hence interest cost savings = 22636 - interest paid on new loan = 22636-11886 = 10749

Total savings cost = Interest savings - total fee = 10749 - 13180 = -2431